# RL3: Boosting Meta Reinforcement Learning via RL inside RL2

### v5

@@ Abstract
Meta reinforcement learning (meta-RL) methods such as RL2 have emerged as promising approaches for learning data-efficient RL algorithms tailored to a given task distribution. However, they show poor asymptotic performance and struggle with out-of-distribution tasks because they rely on sequence models, such as recurrent neural networks or transformers, to process experiences rather than summarize them using general-purpose RL components such as value functions. In contrast, traditional RL algorithms are data-inefficient as they do not use domain knowledge, but do converge to an optimal policy in the limit. We propose RL3, a principled hybrid approach that incorporates action-values, learned per task via traditional RL, in the inputs to meta-RL. We show that RL3 earns greater cumulative reward in the long term compared to RL2 while drastically reducing meta-training time and generalizes better to out-of-distribution tasks. Experiments are conducted on both custom and benchmark discrete domains from the meta-RL literature that exhibit a range of short-term, long-term, and complex dependencies.
@@ 1 Introduction
Reinforcement learning (RL) has been shown to produce effective policies in a variety of applications including both virtual [1] and embodied [2, 3] systems. However, traditional RL algorithms have three major drawbacks: they can be slow to converge, require a large amount of data, and often have difficulty generalizing to out-of-distribution (OOD) tasks not practiced during training. These shortcomings are especially glaring in settings where the goal is to learn policies for a collection or distribution of problems that share some similarities, and for which traditional RL must start from scratch for each problem. For example, many robotic manipulation tasks require interacting with an array of objects with similar but not identical shapes, sizes, weights, materials, and appearances, such as mugs and cups. It is likely that effective manipulation strategies for these tasks will be similar, but they may also differ in ways that make it challenging to learn a single policy that is highly successful on all instances. Recently, meta reinforcement learning (meta-RL) has been proposed as an approach to mitigate these shortcomings by deriving RL algorithms (or meta-RL policies) that adapt efficiently to a distribution of tasks that share some common structure [4, 5].
While meta-RL systems represent a significant improvement over traditional RL in such settings, they still require large amounts of data during meta-training time, can have poor asymptotic performance during adaptation, and although they “learn to learn," they often generalize poorly to tasks not represented in the meta-training distribution. This is partly because they rely on black-box sequence models like recurrent neural networks or transformers to process experience data. These models cannot handle arbitrary amounts of data effectively and lack integrated general-purpose RL components that could induce a broader generalization bias.
@@
While Q-value estimates can be injected into any other meta-RL algorithm, for clarity of exposition, we implement RL3 by injecting Q-value estimates into one of the most popular and easily understood meta-RL algorithm, RL2 [4] (hence, the name RL3). However, it should be noted that our baseline implementation of RL2 includes significant enhancements like using transformers instead of LSTMs to improve long-context reasoning, in addition to incorporating numerous recommendations from Ni et al. [6] that have been shown to make recurrent model-free RL algorithms like RL2 competitive with state-of-the-art meta-RL baselines like VariBAD [7].
The primary contribution of this paper is a proof-of-concept that injecting Q-estimates obtained via traditional object-level RL alongside the typical experience histories within a meta-RL agent leads to higher long-term returns and better OOD generalization, while maintaining short-term efficiency. We further demonstrate that our approach can also work with an abstract, or coarse, representation of the object-level MDP. We experiment with discrete domains that both reflect the challenges faced by meta-RL and simultaneously allow transparent analysis of the results. Finally, we examine the key insights that inform our approach and show theoretically that object-level Q-values are directly related to the optimal meta-value function.
@@ 2 Related Work
Although meta-RL is a fairly new topic of research, the general concept of meta-learning is decades old [8], which, coupled with a significant number of design decisions for meta-RL systems, has created a large number of different proposals for how systems ought to best exploit the resources available within their deployment contexts [9]. At a high level, most meta-RL algorithms can be categorized as either parameterized policy gradient (PPG) models [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] or black box models [4, 23, 5, 24, 25, 26, 27, 28, 7, 29, 30, 31]. PPG approaches assume that the underlying learning process is best represented as a policy gradient, where the set of parameters that define the underlying algorithm ultimately form a differentiable set of meta-parameters that the meta-RL system may learn to adjust. The additional structure provided by this assumption, combined with the generality of policy gradient methods, means that typically PPG methods retain greater generalization capabilities on out-of-distribution tasks. However, due to their inherent data requirements, PPG methods are often slower to adapt and initially train.
@@
In this paper we focus on black box models, which represent the meta-learning function as a neural network, often a recurrent neural network (RNN) [4, 23, 5, 26, 27, 28, 7, 29] or a transformer [25, 32, 33]. There are also several hybrid approaches that combine PPG and black box methods, either during meta-training [34] or fine-tuning [35, 36]. Using black box models simplifies the process of augmenting meta states with Q-estimates and allows us to retain relatively better data efficiency while relying on the Q-value injections for better long-term performance and generalization.
@@
Meta-RL systems may also leverage extra information available during training, such as task identification [26, 29]. Such ‘privileged information’ can of course lead to more performant systems, but is not universally available. As our hypothesis does not rely on the availability of such information, we expect our approach to be orthogonal to, and compatible with, such methods. Black box meta-RL systems that do not use privileged information still vary in several ways, including the choice between on-policy and off-policy learning and, in systems that use neural networks, the choice between transformers [37] and RNNs [38, 39, 40].
@@
The most relevant methods to our work are end-to-end methods, which use a single function approximator to subsume both learner and meta-learner, such as RL2 [4], L2L [5], SNAIL [25], and E-RL2 [16], and methods that exploit the formal description of the meta-RL problem as a POMDP or a Bayes-adaptive MDP (BAMDP) [41]. These methods attempt to learn policies conditioned on the BAMDP belief state while also approximating this belief state by, for example, variational inference (VariBAD) [7, 42], or random network distillation on belief states (HyperX) [43]. Or, they simply encode enough experience history to approximate POMDP beliefs (RL2) [4, 5].
@@ 3.1 Partially Observable MDPs
We use the standard notation defining a Markov decision process (MDP) as a tuple $M=\langle S,A,T,R\rangle$, where $S$ is a set of states; $A$ is a set of actions; $T$ is the transition and $R$ is the reward function. A partially observable Markov decision process (POMDP) extends MDPs to settings with partially observable states. A POMDP is described as a tuple $\langle S,A,T,R,\Omega,O\rangle$, where $S,A,T,R$ are as in an MDP. $\Omega$ is the set of possible observations, and $O:S\times A\times\Omega\rightarrow[0,1]$ is an observation function representing the probability of receiving observation $\omega$ after performing action $a$ and transitioning to state $s^{\prime}$. POMDPs can alternatively be represented as continuous-state belief-MDPs where a belief state $b\in\Delta^{|S|}$ is a probability distribution over all states. In this representation, a policy $\pi$ is a mapping from belief states to actions, $\pi:\Delta^{|S|}\rightarrow A$.
@@ 3.2 Reinforcement Learning
Reinforcement learning (RL) agents learn an optimal policy given an MDP with unknown dynamics using only transition and reward feedback. This is often done by incrementally estimating the optimal action-value function $Q^{*}(s,a)$ [44], which satisfies the Bellman optimality equation $Q^{*}(s,a)=\mathbb{E}_{s^{\prime}}[R(s,a)+\gamma\max_{a^{\prime}\in A}Q^{*}(s^{\prime},a^{\prime})]$. In large or continuous state settings, it is popular to use deep neural networks to represent the action-value functions [1]. We denote the vector representing the Q-estimates of all actions at state $s$ as $Q(s)$, and after $t$ feedback steps, as $Q^{t}(s)$. Q-learning is known to converge asymptotically [45], provided each state-action pair is explored sufficiently. As a rough general statement, $||Q^{t}(s)-Q^{*}(s)||_{\infty}$ is proportional to $\approx\frac{1}{\sqrt{t}}$, with strong results on the convergence error available [46, 47, 48]. The theoretical objective in RL is to optimize the value of the final policy i.e., the cumulative reward per episode, disregarding the data cost incurred and the cumulative reward missed (or regret) during learning due to suboptimal exploration.
@@ 3.3 Meta Reinforcement Learning
Meta-RL seeks action selection strategies that minimize regret in MDPs drawn from a distribution of MDPs that share the same state and action spaces. Therefore, the objective in meta-RL is to maximize the cumulative reward over the entire interaction (or adaptation) period with an MDP, which may span multiple episodes, in order to optimize the exploration-exploitation tradeoff. Formally,
$\mathcal{J}(\theta)=\mathbb{E}_{M_{i}\sim\mathcal{M}}\Big{[}\sum_{t=0}^{H}\gamma^{t}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{i}^{\pi_{\theta}}}[R_{i}(s_{t},a_{t})]\Big{]}$ (1)
where the meta-RL policy $\pi_{\theta}$ is interpreted as a ‘fast’ or ‘inner’ RL algorithm that maps the experience sequence $(s_{0},a_{0},r_{0},...,s_{t})$ within an MDP $M_{i}$ to an action $a_{t}$ using either a recurrent neural network or a transformer network. $\rho_{i}^{\pi_{\theta}}$ is the state-action occupancy induced by the meta-RL policy in MDP $M_{i}$, and $H$ is the length of the adaptation period, or interaction budget. The objective $\mathcal{J}(\theta)$ is maximized using a conventional ‘slow’ or ‘outer’ deep RL algorithm, given the reformulation of the interaction period with an MDP as a single (meta-)episode in the objective function, which maximizes the cumulative reward throughout this period. We will use the term ‘experience history’, denoted by $\Upsilon$, to refer to the state-action-reward sequence within a meta-episode, which spans across multiple episodes $\{\tau_{0},\tau_{1},...\tau_{n}\}$. Fig. 1 illustrates how these components interconnect.
Another way to conceptualize this problem is to recognize that the meta-RL problem may be written as a meta-level POMDP, where the hidden variable is the particular MDP (or task) at hand, $M_{i}$, which varies across meta-episodes. This framing, known as Bayesian RL [49], leverages the fact that augmenting the task-specific state $s$ with belief over tasks $b(i)$ results in a Markovian meta-state $[s,b]$ for optimal action selection, a model known as the Bayes Adaptive MDP (or BAMDP) [41]. That is, this belief state captures all requisite information for the purpose of acting. We will revisit this concept to develop intuition on the role of object-level Q-value estimates in the meta-RL value function.
@@ 4 RL3
To address the limitations of black box meta-RL methods, we propose RL3, a principled approach that leverages (1) the inherent generality of action-value estimates, (2) their ability to compress experience histories into useful summaries, (3) their direct actionability & asymptotic optimality, (4) their ability to inform task-identification, and (5) their relation to the optimal meta-value function, in order to enhance out-of-distribution (OOD) generalization and performance over extended adaptation periods. The central, novel mechanism in RL3 is an additional ‘object-level’ RL procedure executed within the meta-RL architecture, shown in Fig. 1, that computes task-specific optimal Q-value estimates $Q^{t}_{i}(s_{t})$ and state-action counts as supplementary inputs to the meta-RL policy in conjunction with the sequence of states, actions and rewards $(s_{0},a_{0},r_{0},...,s_{t})$. The Q-estimates are computed off-policy, and may involve model estimation and planning, for greater data efficiency. The estimates and the counts are reset at the beginning of each meta-episode as a new task $M_{i}$ is sampled. In all subsequent text, Q-value estimates used as input entail the inclusion of state-action counts as well. We now present a series of key insights informing our approach.
@@
First, estimating action-values is a key component in many universal RL algorithms, and asymptotically, they fully inform optimal behavior irrespective of domain. Strategies for optimal exploration-exploitation trade-off are domain-dependent and rely on historical data, yet many exploration approaches use estimated Q-values and some notion of counts alone, such as epsilon-greedy, Boltzmann exploration, upper confidence bounds (UCB/UCT) [50, 51], count-based exploration [52], curiosity based exploration [53] and maximum-entropy RL [3]. This creates a strong empirical case that using Q-value estimates and state-action counts for efficient exploration has inherent generality.
@@
Second, Q-estimates summarize experience histories of arbitrary length and order in one constant-size vector. This mapping is many-to-one, and any permutation of transitions ($\langle s,a,r,s^{\prime}\rangle$ tuples) or episodes in a history of experiences yield the same Q-estimates. Although this compression is lossy, it still “remembers" important aspects of the experienced episodes, such as high-return actions and goal positions (see Fig. 2) since Q-estimates persist across episodes. This simplifies the mapping the meta agent needs to learn as Q-estimates represent a smaller and more salient set of inputs compared to all possible histories with the same implication.
@@
Third, Q-estimates are actionable. Estimated off-policy, they explicitly represent the optimal exploitation policy for the current task given the data insofar as the RL module is data-efficient, relieving the meta-RL agent from performing such calculations inside the transformer/RNN. Over time, Q-estimates become more reliable and directly indicate the optimal policy whereas processing raw data becomes more challenging. Fortunately, by incorporating Q-estimates the meta-RL agent can eventually ignore the history in the long run (or towards the end of the interaction period) and simply exploit the Q-estimates by selecting actions greedily.
@@
Fourth, Q-estimates are excellent task discriminators and serve as another line of evidence vis-à-vis maintaining belief over tasks. In a simple domain like Bernoulli multi-armed bandits [4], Q-estimates and action-counts combined are sufficient for Bayes-optimal behavior even without providing raw experience data – a result surprisingly unstated in the literature to the best of our knowledge (see Appendix A.1). However, Q-estimates and action-counts may not always be sufficient for Bayes-optimal beliefs [^1] For example, in Gaussian multi-armed bandits, the sufficient statistics include the variance in rewards for each action (see Appendix A.2).. In more complex domains, it is hard to prove the sufficiency of Q-estimates regarding task discrimination. However, via empirical analysis in Appendix D, we argue that i) it is highly improbable for two tasks to have similar Q functions and ii) Q-estimates tend to become accurate task predictors in just a few steps. This implies that the meta-agent may use this finite summary for task inference rather than relying completely on arbitrarily long histories, potentially contributing to enhanced performance over long adaptation periods.
@@
It can be theoretically argued that since the meta agent is a BAMDP policy, it is meta-trained to select greedy actions w.r.t. the BAMDP meta-value function and thus should not require constructing a task-specific plan internally. However, the optimality of the meta action-value function depends on implicitly (or explicitly in some approaches [26, 7, 42, 43]) maintaining a Bayes-optimal belief over tasks in the transformer/RNN architecture. This may be challenging if the task distribution is too broad and the function approximator is not powerful enough to integrate experience histories into Bayes-optimal beliefs, or altogether impossible if there is a distribution shift at meta-test time. This latter condition is common in practice and is a frequent target use case for meta-RL systems. Incorporating task-specific Q-estimates gives the agent a simple alternative (even if not Bayes-optimal) line of reasoning to translate experiences into actions. Incorporating Q-estimates thus reduces susceptibility to distribution shifts since the arguments presented in this section are domain independent.
@@
Finally, Q-estimates often converge far more quickly than the theoretical rate of $\frac{1}{\sqrt{t}}$, allowing them to be useful in the short and medium term, since i) most real-world domains contain significant determinism, ii) it is not necessary to estimate Q-values for states unreachable by the optimal policy, and iii) optimal meta-RL policies may represent active exploration strategies in which Q-estimates converge faster, or evolve in a manner leading to quicker task identification. This is intuitively apparent in shortest-path problems, as illustrated in Fig. 2(a). In a deep neural network, it is difficult to know exactly how Q-estimates will combine with state-action-reward histories when approximating the meta-value function. However, as we show below, we can write an equation for the meta-value function in terms of these constituent streams of information, which may explain why this function is seemingly relatively easy to learn compared to predicting meta-values from histories alone.

### v6

@@ Abstract
Meta reinforcement learning (Meta-RL) methods such as RL2 have emerged as promising approaches for learning data-efficient RL algorithms tailored to a given task distribution. However, they show poor asymptotic performance and struggle with out-of-distribution tasks because they rely on sequence models, such as recurrent neural networks or transformers, to process experiences rather than summarize them using general-purpose RL components such as value functions. In contrast, traditional RL algorithms are data-inefficient as they do not use domain knowledge, but do converge to an optimal policy in the limit. We propose RL3, a principled hybrid approach that incorporates action-values, learned per task via traditional RL, in the inputs to Meta-RL. We show that RL3 earns a greater cumulative reward in the long term compared to RL2 while drastically reducing meta-training time and generalizes better to out-of-distribution tasks. Experiments are conducted on Meta-RL benchmarks and custom discrete domains that exhibit a range of short-term, long-term, and complex dependencies.
@@ 1 Introduction
Reinforcement learning (RL) has been shown to produce effective policies in a variety of applications, including both virtual (Mnih et al., 2015) and embodied (Schulman et al., 2017; Haarnoja et al., 2018) systems. However, traditional RL algorithms have three major drawbacks: they can be slow to converge, require a large amount of data, and often have difficulty generalizing to out-of-distribution (OOD) tasks not practiced during training. These shortcomings are especially glaring in settings where the goal is to learn policies for a collection or distribution of problems that share some similarities, and for which traditional RL must start from scratch for each problem. For example, many robotic manipulation tasks require interacting with an array of objects with similar but not identical shapes, sizes, weights, materials, and appearances, such as mugs and cups. It is likely that effective manipulation strategies for these tasks will be similar, but they may also differ in ways that make it challenging to learn a single policy that is highly successful on all instances. Recently, meta-reinforcement learning—often dubbed “learning to learn”— has been proposed as an approach to mitigate these shortcomings by deriving RL algorithms (or Meta-RL policies) that adapt efficiently to a distribution of tasks that share some common structure (Duan et al., 2016; Wang et al., 2016).
@@
While Meta-RL systems represent a significant improvement over traditional RL in such settings, they still require large amounts of data during meta-training time, can have poor asymptotic performance during adaptation, and may generalize poorly to tasks not represented in the meta-training distribution. This is partly because they rely on black-box sequence models like recurrent neural networks or transformers to process experience data. These models cannot handle arbitrary amounts of data effectively and lack integrated general-purpose RL components that could induce a broader generalization bias.
@@
While $Q$-value estimates can be injected into any other Meta-RL algorithm, for clarity of exposition, we implement RL3 by injecting $Q$-value estimates into one of the most popular and easily understood Meta-RL algorithm, RL2 (Duan et al., 2016), hence, the name RL3. However, it should be noted that our baseline implementation of RL2 includes significant enhancements like using transformers instead of LSTMs to improve long-context reasoning, in addition to incorporating numerous recommendations from Ni et al. (2022) that have been shown to make even recurrent model-free RL algorithms like RL2 competitive with strong Meta-RL baselines like VariBAD (Zintgraf et al., 2020).
The primary contribution of this paper is a proof-of-concept that injecting $Q$-estimates obtained via traditional object-level RL, alongside experience histories, leads to improved performance, which to some may be surprising as this information is already latent within the original input sequence. Specifically, we observe higher long-term returns and better OOD generalization, while maintaining short-term efficiency in Meta-RL agents like RL2 that run an ‘outer-loop’ deep RL algorithm and use sequence models like LSTMs or transformers. We further demonstrate that our approach can also work with an abstract, or coarse, representation of the object-level MDP. We experiment with discrete domains that both reflect the challenges faced by Meta-RL and simultaneously allow transparent analysis of the results. Finally, we examine the key insights that inform our approach and show theoretically that object-level $Q$-values are directly related to the optimal meta-value function.
@@ 2 Related Work
Although Meta-RL is a fairly new topic of research, the general concept of meta-learning is decades old (Vilalta & Drissi, 2002), which, coupled with a significant number of design decisions for Meta-RL systems, has created a large number of different proposals for how systems ought to best exploit the resources available within their deployment contexts (Beck et al., 2023). At a high level, most Meta-RL algorithms can be categorized as either parameterized policy gradient (PPG) models (Finn et al., 2017; Li et al., 2017; Sung et al., 2017; Al-Shedivat et al., 2018; Gupta et al., 2018; Yoon et al., 2018; Stadie et al., 2018; Vuorio et al., 2019; Zintgraf et al., 2019; Raghu et al., 2019; Kaushik et al., 2020; Ghadirzadeh et al., 2021; Mandi et al., 2022) or black box models (Duan et al., 2016; Heess et al., 2015; Wang et al., 2016; Foerster et al., 2018; Mishra et al., 2018; Humplik et al., 2019; Fakoor et al., 2020; Yan et al., 2020; Zintgraf et al., 2020; Liu et al., 2021; Emukpere et al., 2021; Beck et al., 2022). PPG approaches assume that the underlying learning process is best represented as a policy gradient, where the set of parameters that define the underlying algorithm ultimately form a differentiable set of meta-parameters that the Meta-RL system may learn to adjust. The additional structure provided by this assumption, combined with the generality of policy gradient methods, means that typically PPG methods retain greater generalization capabilities on out-of-distribution tasks. However, due to their inherent data requirements, PPG methods are often slower to adapt and initially train.
@@
In this paper we focus on black box models, which represent the meta-learning function as a neural network, often a recurrent neural network (RNN) (Duan et al., 2016; Heess et al., 2015; Wang et al., 2016; Humplik et al., 2019; Fakoor et al., 2020; Yan et al., 2020; Zintgraf et al., 2020; Liu et al., 2021) or a transformer (Mishra et al., 2018; Wang et al., 2021; Melo, 2022). There are also several hybrid approaches that combine PPG and black box methods, either during meta-training (Ren et al., 2023) or fine-tuning (Lan et al., 2019; Xiong et al., 2021). Using black box models simplifies the process of augmenting meta states with $Q$-estimates and allows us to retain relatively better data efficiency while relying on the $Q$-value injections for better long-term performance and generalization.
@@
Meta-RL systems may also leverage extra information available during training, such as task identification (Humplik et al., 2019; Liu et al., 2021). Such ‘privileged information’ can of course lead to more performant systems, but is not universally available. As our hypothesis does not rely on the availability of such information, we expect our approach to be orthogonal to, and compatible with, such methods. Black box Meta-RL systems that do not use privileged information still vary in several ways, including the choice between on-policy and off-policy learning and, in systems that use neural networks, the choice between transformers (Vaswani et al., 2017) and RNNs (Elman, 1990; Hochreiter & Schmidhuber, 1997; Cho et al., 2014).
@@
The most relevant methods to our work are end-to-end methods, which use a single function approximator to subsume both learner and meta-learner, such as RL2 (Duan et al., 2016), L2L (Wang et al., 2016), SNAIL (Mishra et al., 2018), and E-RL2 (Stadie et al., 2018), and methods that exploit the formal description of the Meta-RL problem as a POMDP or a Bayes-adaptive MDP (BAMDP) (Duff, 2002). These methods attempt to learn policies conditioned on the BAMDP belief state while also approximating this belief state by, for example, variational inference (VariBAD) (Zintgraf et al., 2020; Dorfman et al., 2020), or random network distillation on belief states (HyperX) (Zintgraf et al., 2021). Or, they simply encode enough experience history to approximate POMDP beliefs (RL2) (Duan et al., 2016; Wang et al., 2016).
@@ 3.1 Partially Observable MDPs
We use the standard notation defining a Markov decision process (MDP) as a tuple $M=\langle S,A,T,R\rangle$, where $S$ is a set of states, $A$ is a set of actions, $T$ is the transition function, and $R$ is the reward function. A partially observable Markov decision process (POMDP) extends MDPs to settings with partially observable states. A POMDP is described as a tuple $\langle S,A,T,R,\Omega,O\rangle$, where $S$, $A$, $T$, and $R$ are as in an MDP. In a POMDP, the agent receives observations $\omega\in\Omega$ instead of direct state information, with an observation function $O(\omega|s^{\prime},a)$ representing the probability of observing $\omega$ after taking action $a$ and transitioning to state $s^{\prime}$. POMDPs can also be formulated as continuous-state belief MDPs, where a belief state $b\in\Delta^{|\mathcal{S}|}$ is a probability distribution over all possible states. Beliefs are updated online via Bayesian inference: $b^{\prime}(s^{\prime}|b,a,\omega)\propto O(\omega|s^{\prime},a)\,\sum_{s\in S}T(s^{\prime}|s,a)\,b(s)$ as new observations are received. The policy maps beliefs to actions, $\pi:\Delta^{|\mathcal{S}|}\rightarrow\mathcal{A}$.
@@ 3.2 Reinforcement Learning
Reinforcement learning (RL) agents learn an optimal policy in an MDP with unknown dynamics using only transition and reward feedback. This is often done by incrementally estimating the optimal action-value function $Q^{*}(s,a)$ (Watkins & Dayan, 1992), which satisfies the Bellman optimality equation $Q^{*}(s,a)=\mathbb{E}_{s^{\prime}}[R(s,a)+\gamma\max_{a^{\prime}\in A}Q^{*}(s^{\prime},a^{\prime})]$. In large or continuous state spaces, it is common to use deep neural networks to represent the action-value function (Mnih et al., 2015). We denote the vector of $Q$-estimates for all actions at state $s$ as $Q(s)$, and after $t$ feedback steps as $Q^{t}(s)$. Q-learning is known to converge asymptotically (Sutton & Barto, 2018), provided each state-action pair is explored sufficiently. As a rough statement, $||Q^{t}(s)-Q^{*}(s)||_{\infty}$ is proportional to $\approx\frac{1}{\sqrt{t}}$, with strong results on convergence error available (Szepesvári, 1997; Kearns & Singh, 1998; Even-Dar et al., 2003). Typically, the formal objective in RL is to optimize the value of the final policy, i.e., the cumulative reward per episode, disregarding the cumulative reward missed (or regret) during learning due to suboptimal actions taken by the behavior policy used for exploration.
@@ 3.3 Meta Reinforcement Learning
While standard model-free RL seeks a single optimal policy per MDP, Meta-RL aims to train strategies that rapidly adapt across a distribution of related tasks $\mathcal{M}$. This is especially valuable when the agent faces a fixed interaction budget $H$ per task, assuming shared state and action spaces. Formally, Meta-RL optimizes a ‘fast’ behavior policy $\pi$, also called the Meta-RL policy, to maximize cumulative reward (or equivalently, minimize regret) during the adaptation period:
$\mathcal{J}(\pi)=\mathbb{E}_{M_{i}\sim\mathcal{M}}\Bigl{[}\sum_{t=0}^{H}\gamma^{t}\,\mathbb{E}_{(s_{t},a_{t})\sim\rho_{i,t}^{\pi}}\bigl{[}R_{i}(s_{t},a_{t})\bigr{]}\Bigr{]},$ (1)
where $\rho_{i,t}^{\pi}$ denotes the state-action distribution induced by $\pi$ in $M_{i}$ at timestep $t$. Unlike Markovian policies in standard RL, the Meta-RL policy conditions on the experience history $\Upsilon=(s_{0},a_{0},r_{0},\dots,s_{t})$, which spans multiple episodes within each adaptation period.
[…]
@@ 4 RL3
To address the limitations of black box Meta-RL methods, we propose RL3, a principled approach that leverages (1) the inherent generality of action-value estimates, (2) their ability to compress experience histories into useful summaries, (3) their direct actionability & asymptotic optimality, (4) their ability to inform task identification, and (5) their relation to the optimal meta-value function, in order to enhance out-of-distribution (OOD) generalization and performance over extended adaptation periods. The central, novel mechanism in RL3 is an additional ‘object-level’ RL procedure executed within the Meta-RL architecture, shown in Fig. 1, that computes task-specific optimal $Q$-value estimates $Q^{t}_{i}(s_{t})$—denoting $\hat{Q}^{*}$ estimates for task $i$ after $t$ transition samples—and state-action counts as supplementary inputs to the Meta-RL policy in conjunction with the sequence of states, actions, and rewards $(s_{0},a_{0},r_{0},...,s_{t})$. The $Q$-estimates are computed off-policy and may involve experience replay or model estimation & planning for greater data efficiency. The estimates and the counts are reset at the beginning of each meta-episode as a new task $M_{i}$ is sampled. We now present a series of key insights informing our approach.
@@
First, estimating action-values is a key component in many general-purpose, or universal, RL algorithms, and asymptotically, they fully inform optimal behavior irrespective of domain. Strategies for optimal exploration-exploitation trade-off are domain-dependent and rely on historical data, yet many exploration approaches use estimated $Q$-values and some notion of counts alone, such as epsilon-greedy, Boltzmann exploration, upper confidence bounds (UCB/UCT) (Auer, 2002; Kocsis & Szepesvári, 2006), count-based exploration (Tang et al., 2017), curiosity-based exploration (Burda et al., 2019), and maximum-entropy RL (Haarnoja et al., 2018). This creates a strong empirical case that using $Q$-value estimates and state-action counts for efficient exploration has inherent generality.
@@
Second, $Q$-estimates summarize experience histories of arbitrary length and order in one constant-size vector. This mapping is many-to-one, and any permutation of transitions ($\langle s,a,r,s^{\prime}\rangle$ tuples) or episodes in a history of experiences yield the same $Q$-estimates. Although this compression is lossy, it still “remembers” important aspects of the experienced episodes, such as high-return actions and goal positions (see Fig. 2), since $Q$-estimates persist across episodes. In conjunction, action counts provide an explicit record of exploration coverage. This simplifies the mapping the meta-agent needs to learn, as $Q$-estimates and counts together represent a smaller and more salient set of inputs compared to all possible histories with the same implication.
@@
Third, $Q$-estimates are actionable. Estimated off-policy, they explicitly represent the optimal exploitation policy for the current task given the data, insofar as the RL module is data-efficient, relieving the Meta-RL agent from performing such calculations inside the transformer or RNN. Over time, $Q$-estimates become more reliable and directly indicate the optimal policy, whereas processing raw data becomes more challenging. Provided alongside action counts, they also give the agent an implicit mechanism to modulate trust in $Q$-estimates as evidence accumulates. Fortunately, by incorporating $Q$-estimates, the Meta-RL agent can eventually ignore the history in the long run (or towards the end of the interaction period) and simply exploit the $Q$-estimates by selecting actions greedily.
@@
Fourth, $Q$-estimates are excellent task discriminators and serve as another line of evidence vis-à-vis maintaining belief over tasks. In a simple domain like Bernoulli multi-armed bandits (Duan et al., 2016), $Q$-estimates and action counts combined are sufficient for Bayes-optimal behavior even without providing raw experience data—a result surprisingly unstated in the literature to the best of our knowledge (see Appendix A.1). However, $Q$-estimates and counts may not always be sufficient for Bayes-optimal beliefs (see Gaussian multi-armed bandits example in Appendix A.2). In more complex domains, it is hard to prove the sufficiency of $Q$-estimates regarding task discrimination. However, via empirical analysis in Appendix D, we argue that i) it is highly improbable for two tasks to have similar Q functions and ii) $Q$-estimates tend to become accurate task predictors in just a few steps. This implies that the meta-agent may use this finite summary for task inference rather than relying completely on arbitrarily long histories, potentially contributing to enhanced performance over long adaptation periods.
@@
It can be theoretically argued that since the meta-agent is a BAMDP policy, it is meta-trained to select greedy actions with respect to the BAMDP meta-value function and thus should not require constructing a task-specific plan internally. However, the optimality of the meta action-value function depends on implicitly (or explicitly in some approaches (Humplik et al., 2019; Zintgraf et al., 2020; Dorfman et al., 2020; Zintgraf et al., 2021)) maintaining a Bayes-optimal belief over tasks in the transformer or RNN architecture. This may be challenging if the task distribution is too broad and the function approximator is not powerful enough to integrate experience histories into Bayes-optimal beliefs, or altogether impossible if there is a distribution shift at meta-test time. This latter condition is common in practice and is a frequent target use case for Meta-RL systems. Incorporating task-specific $Q$-estimates gives the agent a simple alternative (even if not Bayes-optimal) line of reasoning to translate experiences into actions. Incorporating $Q$-estimates thus reduces susceptibility to distribution shifts, since the arguments presented in this section are domain independent.
@@
Finally, $Q$-estimates often converge far more quickly than the theoretical rate of $1/\sqrt{t}$, allowing them to be useful in the short and medium term, since i) most real-world domains contain significant determinism, ii) it is not necessary to estimate $Q$-values for states unreachable by the optimal policy, and iii) optimal Meta-RL policies may represent active exploration strategies in which $Q$-estimates converge faster, or evolve in a manner leading to quicker task identification. This is intuitively apparent in shortest-path problems, as illustrated in Fig. 2(a). In a deep neural network, it is difficult to know exactly how $Q$-estimates will combine with state-action-reward histories when approximating the meta-value function. However, as we show below, we can write an equation for the meta-value function in terms of these constituent streams of information, which may explain why this function is seemingly relatively easy to learn compared to predicting meta-values from histories alone.